# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v1

@@ 1 Introduction
English, as the global lingua franca, is predominant in large-scale text corpora used to train Large Language Models (LLMs) (Ziems et al., 2023; Zhang et al., 2023), including widely used datasets like CommonCrawl. These datasets are primarily tailored to an English-speaking audience located in the United States, and are mainly composed of privileged English dialects from wealthier educated urban zones (Talat et al., 2022; Ziems et al., 2023; Ryan et al., 2024; Gururangan et al., 2022). This biased training dataset composition permeates the LLM, resulting in models tailored to these English dialects (Santy et al., 2023; Hall et al., 2022). This highlights underlying design biases in LLMs, a phenomenon where certain design choices result in improved downstream performance for specific sub-populations (Santy et al., 2023). Consequently, their effectiveness considerably decreases when prompted in other languages or even in underrepresented English dialects (Lai et al., 2023; Zhang et al., 2023; Bang et al., 2023; Ziems et al., 2023; Ryan et al., 2024).
@@
Figure 1: An example prompt of a native and non-native English speaker and the corresponding output given by GPT4o. The desired output is C. The model selects an incorrect answer choice for the non-native English speaker, although semantically the same message was conveyed.
@@
LLMs are highly sensitive to prompt formulations Beck et al. (2024); Chakraborty et al. (2023).  Ryan et al. (2024) show how models’ responses are tailored to Western English dialects, with prompt selection impacting LLMs’ preference tuning. Therefore, prompting models in other dialects can result in performance differences due to these design biases. Ziems et al. (2023) even provide a dataset covering multiple English dialects. However, unlike those studies focusing only on English dialects from English-speaking countries, our research also incorporates participants from countries where English is not an official language.
@@
In this paper, we find performance differences when LLMs are prompted by native versus non-native English speakers. More specifically, models often generate inaccurate responses for non-native speakers and rate the native prompts more positively than intended. We collect a dataset comprising over 12,000 unique prompts from native and non-native English speakers around the world, and demonstrate how different prompt formulations can lead to worse performance despite conveying the same message. An example of a prompt in our dataset is shown in Figure 1. Moreover, we find that these performance differences increase when comparing native English speakers from Western countries (US, UK, Canada) with other native and non-native English speakers. Furthermore, when the model recognizes or is informed about the user’s nativeness, a strong anchoring effect occurs, where the added information substantially affects model performance, leading to increased bias towards native English speakers.
Our contributions are as follows:
  • We quantitatively and qualitatively analyze the performance of LLMs on objective and subjective classification tasks, as well as generative tasks, when prompted by native and non-native English speakers
  • We investigate the impact on the LLM performance when the nativeness of the user is explicitly stated or inferred by the model.
@@
  • We publish our multilingual instruction-tuning dataset [^1] https://github.com/manon-reusens/native_en_bias containing over 12,000 unique prompts from a diverse group of native and non-native English speakers worldwide, including translations of the prompts into eight different native languages.
@@ 2 Related work
Model Positionality and Design Bias. Model positionality, coined by Cambo and Gergle (2022), refers to the social and cultural position of a model, influenced by the stakeholders involved in its development, such as annotators and developers. This positionality affects the inclusivity of LLMs, as they evolve with certain biases that may disadvantage specific populations. (Cambo and Gergle, 2022; Santy et al., 2023). Design biases arise when researchers make choices that improve model performance for specific sub-populations Santy et al. (2023). A notable example is the overrepresentation of English pretraining corpora, which leads to disproportionate performance improvements in English compared to other languages (Qin et al., 2023; Blasi et al., 2022; Joshi et al., 2020).
@@ 3 Methodology
Given the sensitivity of LLMs to prompt formulation (Beck et al., 2024; Chakraborty et al., 2023), the diversity of English dialects (Ziems et al., 2023; Ryan et al., 2024), and alignment of models towards Western native English speakers (Ryan et al., 2024; Santy et al., 2023; Gururangan et al., 2022), we hypothesize that these design choices affect LLM performance when interacting with native versus non-native English speakers. Furthermore, we hypothesize that the specific English dialect learned may influence the responses from LLMs, particularly if the LLMs are optimized for English variants from the US, UK, and other predominantly English-speaking countries, as shown by i.e. Santy et al. (2023). We also anticipate that performance differences will increase when a model is explicitly informed about the user’s nativeness. Considering the LLM’s ability to infer demographic annotator information from text (Staab et al., 2023), we also expect the LLM to infer a user’s nativeness and that this correct inference leads to greater performance differences.
@@ 3.1 Dataset
From each original dataset, we randomly selected 100 examples, ensuring they were correctly annotated and free of offensive language. Additionally, we included one extra example per dataset to serve as a tutorial for the annotator to get used to the task.
@@ 3.2 Annotations
We required the annotators to have a minimum level of English equivalent to a high school or university degree to ensure English proficiency. Each annotator worked on 20 to 240 examples, and we gathered them through direct recruitment, opting for an open annotation process rather than an existing annotation platform to ensure high-quality annotations. All annotators were reimbursed at a rate of at least 12.11 euros per hour.
@@
Annotators performed different tasks depending on the assigned datasets. An example annotation is shown in Figure 2, where a task definition is provided together with an impractical statement. The annotator has to provide the [Annotator PROMPT] based on the task definition and the desired output, which is C in this example. More details about the annotation setup can be found in Appendix B.
Before including the annotations in our final dataset, they were validated. An annotation was deemed invalid if it met any of the following criteria: 1) The response was unrelated to the task, i.e. "I don’t know / understand", or a response for a different topic or question. 2) The response contained (part of) the answer. 3) The response did not follow the required format or task definition. 4) The annotator misunderstood the task. Examples for each validation criterion are included in Appendix C.
@@ 4.1 Gathering LLM responses
Using gathered annotations, we conducted experiments with the chat-versions of well-established LLMs, as these are used in daily life. An overview of the checkpoints per model is shown in Appendix F. We included GPT3.5444https://openai.com/index/gpt-3-5-turbo-fine-tuning-and-api-updates/, GPT4o [^5] https://openai.com/index/hello-gpt-4o/, Haiku (Anthropic, 2024), Sonnet (Anthropic, 2024), using the appropriate APIs, and Qwen1.5 7B [^6] We ran the experiments for Qwen using A100 GPUs. (Bai et al., 2023) in line with the provided licenses and all consistent with the intended use. This set includes models of varying sizes, different performances, and from different developers, ensuring a diverse representation. Moreover, Qwen, developed by Chinese researchers, provides an interesting comparison in terms of design bias.
@@ 4.2 Evaluation
$\Delta_{\text{native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right),\psi\right)$
$\Delta_{\text{non-native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right),\psi\right)$
@@
The performance metrics for generative tasks are calculated using the LLM-generated gold answer as reference. This gold answer was generated using the original prompt in the dataset as displayed in Figure 2. We used this instead of the original gold answer in the dataset, given that optimal answers are model and context-dependent. By employing the LLM-generated gold answers, we align responses for each model. We include the results, using the original gold output in Appendix.
@@ 5 Results
In line with our hypotheses, the results indicate performance differences between native and non-native English speakers and Western native and not Western native speakers for GPT4o, Sonnet, and Qwen for the classification tasks, as shown in Figures 3 and  4. However, the generative tasks indicate no or even opposite preferences as displayed in Figure 5. Furthermore, informing the models about the user’s nativeness further reduces answer quality for non-native English speakers, also resulting in a strong anchoring effect.
@@ 5.1 Classification tasks
Do LLMs perform better for certain groups on objective classification tasks? Figure 3 shows the overall average performance for the different models and groups on the objective classification tasks. These contain all classification tasks, except the Amazon Food review task.
(a) Native vs. non-native
(b) Western native vs. not
Figure 3: Overall average accuracy of the objective classification tasks per model and group.
Figure 3(a) shows that only GPT4o and Qwen 7B perform better for the native group of annotators. All other models provide on-par results for both groups. Interestingly, Qwen, which is predominantly trained on Chinese text, also prefers native English speakers over non-native English speakers. Given that GPT4o is the best-performing model from our set, the results indicate that its superior performance may come at the cost of being more specifically tailored to native English speakers.
Figure 3(b) indicates that beside GPT4o and Qwen 7B, also Sonnet shows a higher performance for the Western native English speakers. Interestingly, the performance difference between the two groups for Qwen has decreased, compared to Figure 3(a), indicating that the model is less tailored towards Western native English.
Similarly, GPT 3.5 performs best for the not Western English speakers. The results suggest that GPT4o, a larger and generally more effective model than GPT3.5, is more sensitive to prompts from native and non-native speakers, leading to the observed performance disparity. Additionally, this behavior is not exclusive to the two GPT models. A comparison between Haiku and Sonnet, both part of the Claude3 family, reveals that Sonnet also exhibits greater sensitivity to prompts from Western native English speakers.
Can we make similar conclusions for the subjective classification task?
@@
Surprisingly, Figure 4 shows the opposite effect in comparison to the objective classification tasks, preferring both the non-native and not Western native English speakers. This finding is remarkable, given the results in the subjective classification literature (Santy et al., 2023; Durmus et al., 2023).
@@
To understand this contrasting phenomenon, we first analyzed the results for both the native and non-native English speakers, and the Western native and not Western native English speakers, in depth. We find that for the native English-speaking group, the models often predict the rating more positively than actually intended, and this effect is even more pronounced for the Western native English speakers. While for the non-native English group and not Western native English speaking group, GPT4o predicted 50% of all wrongly predicted annotations to be more positive than intended, this was around 60% for the native English speaking group and even 70% for the Western native English speaking group for GPT4o indicating cultural differences. More details about these different distributions for the different models can be found in Appendix H.
Although we find that GPT4o seems to be most tailored towards the native English speakers, it tends to estimate native or Western native responses more positive than non-native responses. This implies cultural differences in this subjective classification, indicating that these models do not always align with native or Western native opinions as previously shown in the literature.
(a) Native vs. non-native
(b) Western native vs. not
Figure 4: Overall average accuracy of the subjective classification task per model and group.
What is the effect on the performance when informing the model about the nativeness of the annotator?
@@
For these experiments, we added information about the user’s nativeness to the system prompt. More specifically, we asked the model to respond as if it is interacting with a native or non-native English speaker respectively.
correct information wrong information
model native non-native native non-native
GPT3.5 0.84 0.82 0.84 0.84
GPT4o 0.94 0.93 0.94 0.93
Haiku 0.86 0.84 0.85 0.86
Sonnet 0.86 0.66 0.66 0.86
qwen 7B 0.83 0.82 0.83 0.83
Table 1: Overall average accuracy of all classification tasks per model when adding correct/wrong information about whether or not the prompt was written by a native English speaker.
@@
Table 1 shows that when adding information about the nativeness of the annotator to the system prompt when generating the predictions, the bias towards the native English-speaking group becomes more pronounced. All models show better performance for the native English speaking group in comparison to the non-native English speakers. The biggest performance difference is seen for Sonnet. When looking into these results, we find that Sonnet started answering several questions languages other than English, such as Spanish, French or Indonesian, when responding as if interacting with non-native speakers. This is remarkable and demonstrates a clear anchoring effect, considering we only instructed the model to respond as if interacting with a non-native English speaker without specifying another language.
When inaccurately labeling native speakers as non-native and vice versa, only the Claude models adhered to this incorrect information, preferring the non-native group. GPT4o still performs better on the native prompts, and both Qwen and GPT3.5 now show equal performance, compared to the slight preference towards the native group in our previous experiment. This indicates that GPT4o is most robust against the additional information on the user’s nativeness, concentrating primarily on the prompt itself. It thus seems to be better hedged against direct bias being added. The same applies to GPT3.5 and Qwen, however to a lesser extent given that adding information that mislabels the non-native English speakers as native English speakers does enhance performance.
@@
model guess native guess non-native guess native correctly guess non-native correctly
GPT3.5 0.78 0.77 0.76 0.73
GPT4o 0.83 0.83 0.83 0.81
Haiku 0.65 0.64 0.53 0.64
Sonnet 0.65 0.65 0.53 0.56
qwen 7B 0.69 0.70 0.61 0.76
Table 2: Overall average accuracy of all classification tasks per model and group when guessed by the model whether the person is a native or non-native speaker, and when it is guessed correctly.
Table 2 shows that for Haiku and GPT3.5, first explicitly letting the model guess about the nativeness of the prompt writer, results in a better score for the prompts that were indicated to be native than the ones that were seen as non-native. GPT4o and Sonnet seem on par for both predicted groups. However, when analyzing the performance of the prompts that were guessed correctly, we do see a large performance difference for both groups for GPT3.5 and GPT4o. Haiku, Sonnet and Qwen, on the other hand, show a preference for non-native English speakers. The performance difference for Haiku between the two correctly guessed groups is remarkable. It seems that the model has predicted many well-performing non-native annotations to be actually native. Thus, despite this opposite performance difference, it does show the underlying bias within Haiku to associate better-performing prompts with native speakers. Interestingly, Qwen performs best for the non-native group. For Sonnet, the overall accuracy has dropped significantly in this experimental set-up. Moreover, both Qwen and Sonnet have correctly guessed a very low amount of annotations. This is mainly due to the models’ indecisiveness.

### v2

@@ 1 Introduction
English, as the global lingua franca, is predominant in large-scale text corpora used to train Large Language Models (LLMs) (Ziems et al., 2023; Zhang et al., 2023), including widely used datasets like CommonCrawl. These datasets are primarily tailored to an English-speaking audience located in the United States, and are mainly composed of privileged English dialects from wealthier educated urban zones (Talat et al., 2022; Ziems et al., 2023; Ryan et al., 2024; Gururangan et al., 2022). This biased training dataset composition permeates the LLM, resulting in models tailored to these English dialects (Santy et al., 2023; Hall et al., 2022). This highlights underlying design biases in LLMs, a phenomenon where design choices result in improved downstream performance for specific sub-populations (Santy et al., 2023). Consequently, their effectiveness considerably decreases when prompted in other languages or in underrepresented English dialects (Lai et al., 2023; Zhang et al., 2023; Bang et al., 2023; Ziems et al., 2023; Ryan et al., 2024).
@@
Figure 1: An example prompt of a native and non-native English speaker and the corresponding output given by GPT4o. The desired output is C. The model selects the wrong answer for the non-native English speaker, while semantically the same message was conveyed.
@@
LLMs are highly sensitive to prompt formulations Beck et al. (2024); Chakraborty et al. (2023).  Ryan et al. (2024) show how models’ responses are tailored to Western English dialects, with prompt selection impacting LLMs’ preference tuning. Therefore, prompting models in other dialects can result in performance differences due to these design biases. Ziems et al. (2023) even provide a dataset covering multiple English dialects. However, unlike those studies focusing only on English dialects from English-speaking countries, our research also incorporates participants from countries where English is not an official language. We assess if word sensitivity in prompts disproportionately benefits native English speakers, leading to better model performance. In this case, the model has an inherent native language bias.
@@
In this paper, we find performance differences when LLMs are prompted by native versus non-native English speakers. More specifically, models often generate inaccurate responses for non-native speakers and rate the native prompts more positively than intended. We collect a dataset comprising over 12,000 unique prompts from native and non-native English speakers around the world, and demonstrate how different prompt formulations can lead to worse performance despite conveying the same message. An example of a prompt in our dataset is shown in Figure 1. Moreover, we find that these performance differences increase when comparing native English speakers from Western countries (US, UK, Canada) with other native and non-native English speakers. When the model recognizes or is informed about the user’s nativeness, a strong anchoring effect occurs, where model performance is substantially affected, leading to increased bias towards native English speakers.
@@
Our contributions are as follows: 1) We quantitatively and qualitatively analyze the performance of LLMs on objective and subjective classification tasks [^1] By subjective tasks, we mean classification tasks where the correct answer depends on the subjective interpretation as explained in Beck et al. (2024), as well as generative tasks, when prompted by native and non-native English speakers. 2) We investigate the impact on the LLM performance when the nativeness of the user is explicitly stated or inferred by the model. 3) We publish our multilingual instruction-tuning dataset and code used for the experiments [^2] https://github.com/manon-reusens/native_en_bias containing over 12,000 unique prompts from a diverse group of native and non-native English speakers worldwide, including translations of the prompts into eight different native languages.
@@ 2 Related work
Model Positionality and Design Bias. Model positionality, coined by Cambo and Gergle (2022), refers to the social and cultural position of a model, influenced by the stakeholders involved in its development, such as annotators and developers. This positionality affects the inclusivity of LLMs, as they evolve with certain biases that may disadvantage specific populations (Cambo and Gergle, 2022; Santy et al., 2023). Design biases arise when researchers make choices that improve model performance for specific sub-populations Santy et al. (2023). A notable example is the overrepresentation of English pretraining corpora, which leads to disproportionate performance improvements in English compared to other languages (Qin et al., 2023; Blasi et al., 2022; Joshi et al., 2020).
@@ 3 Methodology
Given the sensitivity of LLMs to prompt formulation (Beck et al., 2024; Chakraborty et al., 2023), the diversity of English dialects (Ziems et al., 2023; Ryan et al., 2024), and alignment of models towards Western native English speakers (Ryan et al., 2024; Santy et al., 2023; Gururangan et al., 2022), we hypothesize that these design choices affect LLM performance when interacting with native versus non-native English speakers. Furthermore, we hypothesize that the specific English dialect learned may influence the responses from LLMs, particularly if the LLMs are optimized for English variants from the US, UK, and other predominantly English-speaking countries, as shown by Santy et al. (2023). We also anticipate that performance differences increase when a model is explicitly informed about the user’s nativeness. Given the LLM’s ability to infer demographic annotator information from text (Staab et al., 2023), we also expect the LLM to infer a user’s nativeness and that this correct inference leads to greater performance differences.
@@ 3.1 Dataset
From each original dataset, we randomly selected 100 examples, manually ensuring they were correctly annotated and free of offensive language. Additionally, one extra example per dataset served as a tutorial for the annotator to get used to the task.
@@ 3.2 Annotations
Annotators had to have a minimum English proficiency level equivalent to a high school or university degree. Each annotator worked on 20 to 240 examples. We gathered them through direct recruitment, opting for an open annotation process rather than an existing annotation platform to ensure high-quality annotations. All annotators were reimbursed at a rate of at least 12.11 euros per hour.
@@
Annotators performed different tasks depending on the assigned datasets. An example annotation is shown in Figure 1, where a task definition is provided together with an impractical statement. The annotator has to provide the [Annotator PROMPT] based on the task definition and the desired output, which is C in this example. We identified the [Annotator PROMPT] per example depending on the dataset. More details about the annotation setup including information about the annotator prompts per dataset can be found in Appendix B.
The authors manually validated the annotations before including them in the final dataset, deeming one invalid if it met any of the following criteria: 1) The response was unrelated to the task, i.e. "I don’t know / understand", or a response for a different topic or question. 2) The response contained (part of) the answer. 3) The response did not follow the required format or task definition. 4) The annotator misunderstood the task. Examples per validation criterion are included in Appendix C.
@@ 4.1 Gathering LLM responses
Using gathered annotations, we conducted experiments with the chat-versions of well-established LLMs, as these are used in daily life. An overview of the checkpoints per model is shown in Appendix F. We included GPT3.5555https://openai.com/index/gpt-3-5-turbo-fine-tuning-and-api-updates/, GPT4o [^6] https://openai.com/index/hello-gpt-4o/, Haiku (Anthropic, 2024), Sonnet (Anthropic, 2024), using the appropriate APIs, and Qwen1.5 7B [^7] We ran the experiments for Qwen using A100 GPUs. (Bai et al., 2023) in line with the provided licenses and all consistent with the intended use. This set includes models of varying sizes, different performances, and from different developers, ensuring a diverse representation. Moreover, Qwen, developed by Chinese researchers, provides an interesting comparison in terms of design bias.
@@ 4.2 Evaluation
$\begin{split}\Delta_{\text{native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{native}}\right),\psi\right)\\ \Delta_{\text{non-native}}=\phi\left(\mathcal{M}\left(\mathcal{T}\mid x_{\text{non-native}}\right),\psi\right)\end{split}$
@@
The performance metrics for generative tasks are calculated using the LLM-generated gold answer as reference. This gold answer was generated using the original prompt in the dataset as displayed in Figure 2. We used this instead of the original gold answer in the dataset, given that optimal answers are model and context-dependent. By employing the LLM-generated gold answers, we align responses for each model.
@@ 5 Results
In line with our hypotheses, the results indicate performance differences between native and non-native English speakers and Western and not Western native speakers for GPT4o, Sonnet, and Qwen for the objective classification tasks, as shown in Table 1. However, the generative tasks indicate an opposite preference as displayed in Table 4. Informing the models about the user’s nativeness further reduces answer quality for non-native English speakers, also resulting in a strong anchoring effect.
@@ 5.1 Classification tasks
Do LLMs perform better for certain groups on both objective and subjective classification tasks? Table 1 shows the average performance and standard deviation over three runs on the objective and subjective classification tasks for the different models and groups. The subjective classification task entails all tasks without an objectively correct answer, as this depends on subjective interpretation. The only subjective task included in our paper is the Amazon Food review task.
objective tasks subjective tasks
native non-native western native not western native native non-native western native not western native
GPT3.5 $0.82\pm 0.01$ $0.81\pm 0.01$ $0.81\pm 0.00$ $0.81\pm 0.01$ $0.76\pm 0.01$ $0.80\pm 0.01$ $0.76\pm 0.00$ $0.80\pm 0.00$
GPT4o $0.94\pm 0.00$ $0.93\pm 0.00$ $0.95\pm 0.01$ $0.93\pm 0.01$ $0.77\pm 0.00$ $0.84\pm 0.01$ $0.77\pm 0.01$ $0.83\pm 0.01$
Haiku $0.87\pm 0.01$ $0.86\pm 0.00$ $0.86\pm 0.00$ $0.87\pm 0.00$ $0.75\pm 0.00$ $0.80\pm 0.00$ $0.71\pm 0.00$ $0.81\pm 0.00$
Sonnet $0.87\pm 0.00$ $0.87\pm 0.00$ $0.87\pm 0.01$ $0.87\pm 0.00$ $0.77\pm 0.00$ $0.79\pm 0.01$ $0.76\pm 0.01$ $0.80\pm 0.00$
Qwen 7B $0.83\pm 0.00$ $0.82\pm 0.01$ $0.83\pm 0.00$ $0.82\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$ $0.72\pm 0.00$ $0.79\pm 0.00$
Table 1: Overall average accuracy of the objective and subjective classification tasks per model and group averaged over three runs and including the standard deviation over these different runs. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
We find that for the objective classification tasks, only GPT4o, Sonnet, and Qwen 7B perform significantly better for the native group of annotators. For the other models, no significant differences are found between the groups. For the western native group, we find that only GPT4o and Qwen 7B a significantly higher performance for the Western native English speakers. Interestingly, we thus conclude that Sonnet is less tailored towards Western native English, which suggests efforts to take into account different English dialects. GPT4o on the other hand, is tailored towards native English speakers, and more specifically to Western native speakers. Interestingly, Qwen, which is predominantly trained on Chinese text, also prefers native and Western native English speakers.
For both the Claude and GPT models, we find that the larger model of the model family is more sensitive to the subtle differences in the prompts of native and non-native speakers. In literature, we find similar positive correlation between model size and bias that slips into models in terms of gender bias Tal et al. (2022). Moreover, Sclar et al. (2023) even show that prompt sensitivity does not decrease for larger model size.
@@
Surprisingly, Table 1 shows the opposite effect for the subjective classification tasks, preferring both the non-native and not Western native English speakers. This finding is remarkable, given the results in the subjective classification literature (Santy et al., 2023; Durmus et al., 2023).
@@
To understand this contrasting phenomenon, we first analyze the results for both the native and non-native English speakers, and the Western native and not Western native English speakers, in depth. We find that for the native English-speaking group, the models often predict the rating more positively than actually intended, and this effect is even more pronounced for the Western native English speakers. While for the non-native English group and not Western native English speaking group, GPT4o predicted 50% of all wrongly predicted annotations to be more positive than intended, this was around 60% for the native English speaking group and even 70% for the Western native English speaking group for GPT4o indicating cultural differences. Appendix H includes the different distributions per model.
@@
What is the effect on the performance when informing the model about the nativeness of the annotator? For these experiments, we added the user’s nativeness to the system prompt. More specifically, we asked the model to respond as if interacting with a (non-)native English speaker.
objective tasks subjective tasks
correct information wrong information correct information wrong information
native non-native native non-native native non-native native non-native
GPT3.5 $0.84\pm 0.00$ $0.81\pm 0.01$ $0.83\pm 0.01$ $0.83\pm 0.01$ $0.76\pm 0.00$ $0.83\pm 0.01$ $0.78\pm 0.00$ $0.83\pm 0.00$
GPT4o $0.94\pm 0.00$ $0.93\pm 0.00$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.77\pm 0.01$ $0.82\pm 0.01$ $0.77\pm 0.01$ $0.84\pm 0.00$
Haiku $0.87\pm 0.01$ $0.84\pm 0.00$ $0.85\pm 0.00$ $0.86\pm 0.00$ $0.75\pm 0.00$ $0.80\pm 0.00$ $0.74\pm 0.00$ $0.81\pm 0.01$
Sonnet $0.88\pm 0.00$ $0.63\pm 0.00$ $0.63\pm 0.00$ $0.87\pm 0.00$ $0.77\pm 0.00$ $0.78\pm 0.00$ $0.78\pm 0.00$ $0.80\pm 0.00$
Qwen 7B $0.83\pm 0.00$ $0.82\pm 0.00$ $0.83\pm 0.00$ $0.83\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$ $0.75\pm 0.00$ $0.79\pm 0.00$
Table 2: Overall average accuracy and standard deviation of the objective and subjective classification tasks per model and group averaged over three runs when adding information about the annotator’s nativeness to the prompt. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold.
@@
Table 2 shows the effect of adding (in)correct information about the nativeness of the annotator to the system prompt on the performance of the models. Compared to Table 1, we now find that all models significantly perform better for the objective classification tasks for the native group compared to the non-native group when correct information is added. When inaccurately labeling native speakers as non-native and vice versa, we see that only GP4o and Qwen still significantly perform better for the native group, suggesting that they are less affected by this additional information. Interestingly, Sonnet is most affected by the information about the (in)correct nativeness of the prompt writer. We find that Sonnet answered several questions in languages other than English, such as Spanish, French or Indonesian, when responding as if interacting with non-native speakers. This is remarkable and shows a clear anchoring effect, considering we only instructed the model to respond as if interacting with a non-native English speaker. Interestingly, this did not occur for the other models
Moreover, also for the other two models that are mostly affected by this anchoring effect, we manually analyzed the errors. GPT3.5 makes more mistakes when informed about the prompt writer being non-native, due to repetition of the instructions, rather than answering the question. Haiku explains the answers, arguing why one option is better than another, thereby failing to follow the instructions, making it hard to assess performance automatically. If both answers are mentioned, we classify the response as inaccurate. Note that this anchoring effect is only seen for the objective classification tasks. It only slightly impacts the subjective classification tasks and does not result in new significant differences than those shown in Table 1. This overall anchoring effect, shows that these three models (GPT3.5, Haiku, and Sonnet) are still sensitive to (useless) information that is given within the prompt as also shown by  Shi et al. (2023); Jones and Steinhardt (2022)
@@
objective tasks subjective tasks
guess native guess non-native guess native correctly guess non-native correctly guess native guess non-native guess native correctly guess non-native correctly
GPT3.5 $0.80\pm 0.01$ $0.77\pm 0.00$ $0.75\pm 0.02$ $0.72\pm 0.01$ $0.75\pm 0.01$ $0.75\pm 0.01$ $0.77\pm 0.03$ $0.77\pm 0.02$
GPT4o $0.85\pm 0.00$ $0.84\pm 0.01$ $0.82\pm 0.05$ $0.82\pm 0.02$ $0.77\pm 0.01$ $0.83\pm 0.01$ $0.74\pm 0.1$ $0.71\pm 0.02$
Haiku $0.65\pm 0.00$ $0.64\pm 0.01$ $0.47\pm 0.03$ $0.66\pm 0.01$ $0.67\pm 0.00$ $0.66\pm 0.01$ $0.74\pm 0.05$ $0.51\pm 0.04$
Sonnet $0.63\pm 0.01$ $0.62\pm 0.01$ $0.52\pm 0.01$ $0.51\pm 0.01$ $0.72\pm 0.01$ $0.77\pm 0.00$ $0.64\pm 0.12$ $0.75\pm 0.04$
Qwen 7B $0.69\pm 0.00$ $0.69\pm 0.00$ $0.74\pm 0.00$ $0.74\pm 0.00$ $0.67\pm 0.00$ $0.71\pm 0.00$ $0.44\pm 0.00$ $0.77\pm 0.00$
Table 3: Overall average accuracy and standard deviation over the three runs of all classification tasks per model and group when guessed by the model whether the person is a native or non-native speaker, and when it is guessed correctly. Significant differences with $p<0.05$ for a Wilcoxon signed-rank test are indicated in bold. No statistical testing is done for the correctly guessed category for the subjective tasks due to the low amount of observations.
Table 3 shows that for GPT3.5, GPT4o, and Sonnet, first explicitly letting the model guess about the nativeness of the prompt writer, results in a significantly better score for the prompts from native than from non-native English speakers, regardless of having correctly guessed the nativeness. Compared to our initial experiments, we thus see that GPT3.5 is more led by this added question and its information and starts performing better for the native class. We find that both GPT models provide accurate explanations of the models’ guesses, recognizing words within the prompt as i.e. British English. Generally, non-native prompts are identified through grammatical errors or awkward phrasing.
The performance difference for Haiku between the two correctly guessed groups is remarkable. The model has predicted many well-performing non-native annotations to be actually native for the objective classification tasks. Qwen now shows a non-significant difference between both. We find that these models struggle with following the guideline to respond with only (non-)native explaining why one is more likely than the other, as does Sonnet. Additionally, Qwen often refrains from guessing, stating it is difficult to determine the nativeness. This results in a lower performance and a non-significant difference between both groups. For the subjective tasks, we find that GPT4o, Sonnet, and Qwen are robust to this prompt, still providing significant differences for the non-native group, while Haiku and GPT3.5 are more impacted.